# More than ‘Left and Right’: Revealing Multilevel Online Political Selective Exposure

### v1

@@ Abstract
Selective exposure, individuals’ inclination to seek out information that supports their beliefs while avoiding information that contradicts them, plays an important role in the emergence of polarization. In the political domain, selective exposure is usually measured on a left-right ideology scale, ignoring finer details. Here, we combine survey and Twitter data collected during the 2022 Brazilian Presidential Election and investigate selective exposure patterns between the survey respondents and political influencers. We analyze the followship network between survey respondents and political influencers and find a multilevel community structure that reveals a hierarchical organization more complex than a simple split between left and right. Moreover, depending on the level we consider, we find different associations between network indices of exposure patterns and 189 individual attributes of the survey respondents. For example, at finer levels, the number of influencer communities a survey respondent follows is associated with several factors, such as demographics, news consumption frequency, and incivility perception. In comparison, only their political ideology is a significant factor at coarser levels. Our work demonstrates that measuring selective exposure at a single level, such as left and right, misses important information necessary to capture this phenomenon correctly.
1 Introduction
@@
The internet, especially social media, has facilitated the online consumption of political information in recent decades [1]. According to a national survey we conducted during the 2022 Brazilian Presidential Election, 82 percent of the Brazilian population acquired political information via social media, with Twitter (now known as X) ranking as one of the most popular social networks for political news. On Twitter, ordinary users mainly receive political information by following certain political influencers, such as politicians, media outlets, journalists, and individual opinion leaders. Despite social networks like Twitter providing a high-choice environment for information consumption and the possibility of deliberative communication, it is a concern for governments and society that users on social networks predominantly follow information aligned with their own opinions, leading to political fragmentation, polarization, and the creation of echo chambers [2, 3, 4, 5, 6]. This phenomenon, known as selective exposure, is attributed to bounded rationality — humans have limited capacity for information processing and prefer to be exposed to information they find agreeable [7] and to the influence of personalized recommendation algorithms on social media platforms [8, 9, 10, 11, 12, 13]. Selective exposure online can, therefore, limit people’s access to diverse opinions and even lead to the spread of extreme views within certain groups [14].
Many studies have explored selective exposure and related topics such as echo chambers, political fragmentation, and polarization on social media, e.g., [15, 16, 17, 18, 19, 20, 5, 21, 22]. In the U.S., for instance, the polarization pattern between Democrats and Republicans has been widely observed by numerous studies such as Refs. [23, 24, 25]. However, in recent years, user-generated content on social media has dramatically altered the dichotomous political dynamics online [26, 27, 28]. It is not only authoritative accounts such as political parties and partisan media outlets that promote political propaganda, but also politicians and individual opinion leaders who create more targeted political information to win elections and achieve business attention [29, 30]. From the perspective of attention economics, if we regard political information as products in the political market, media outlets as the conventional suppliers, and ordinary users as the consumers of political information, the new participation of politicians and other individual opinion leaders has created more substitute products for traditional political information [31]. These substitutes exhibit higher value (more targeted to consumers’ demands) and lower prices (easier interaction functions on social media) compared to conventional suppliers, which leads to a shift of consumers’ attention to these nontraditional suppliers and a redistribution of consumption in the political information market.
@@
In light of this background, political division is more complex in the contemporary era than ever. Evidence from the 2016 U.S. presidential election presented by Bartels et al. [32] highlights that within the Democratic group, various cultural subgroups exist where social identities and other demographic characteristics play a role in forming closely connected communities. Similarly, Brazil’s polarized party systems are also intertwined with complex socio-cultural groups [33]. Particularly, right-wing populists are creating more cleavages among social identity groups, such as gender, ethnic, and religious groups, using social media tools [34]. It is important to understand whether individuals are influenced by these new cleavages and how this affects their consumption of political information online. Our first research question, therefore, aims to understand: what is the complexity of selective exposure among Brazilian online users and how can we capture it? (RQ1). We then continue to investigate which individuals’ attributes are related to selective exposure patterns at systemic levels (RQ2).
@@
To answer the two research questions, we combine survey and Twitter data and connect individual-level attributes and systematic-level patterns. We build connections between consumers and suppliers of political information based on their followingship on Twitter. Consumers are represented by the respondents from a national survey who provided their Twitter handles, while the suppliers are political influencers (including politicians, news media, and individual opinion leaders) followed by consumers on Twitter. Political influencers are annotated based on dimensions of political ideology, campaign support, social identity, and account type. Then, we uncover the complex structures of Brazilian users’ selective exposure on Twitter using a multi-scale community detection approach. The results suggest that ordinary users expose themselves to information based on left-right ideological position, in which granular groups can be found based on other dimensions, such as electoral supports, social identities, and categories of accounts. Therefore, Brazilian users’ selective exposure on Twitter forms a multi-level and hierarchical structure.
@@
We also introduce five indices to measure the selective exposure patterns systematically. By examining a comprehensive set of individuals’ survey attributes with their selective exposure patterns at various levels, we find that, at the lowest level, a wide range of variables, including demographics, ideological position, and news consumption frequency, are significantly associated with the versatility of individuals’ community membership, which we define as Community Overlap. As the level increases, the ideological position becomes more prominent. The other indices are also associated with different individual attributes at various levels.
@@
To the best of our knowledge, this study is the first to explore multi-level selective exposure phenomena in the contemporary online political world. It also makes a significant contribution to the application of the multi-scale community detection approach in social network research, demonstrating that different scales can reveal distinct political ecosystems. The application of multi-scale community detection can enhance our understanding of complex political systems nowadays and shed light on multi-level selective exposure, echo chambers, and polarization online.
1.1 Indices of Selective Exposure Patterns
To better understand the patterns of individuals’ selective exposure behaviors, we develop five indices to measure various patterns of political influencer communities that reflect individuals’ selection. We consider a followship network between political influencers (suppliers) and consumers and define influencer communities as dense groups of influencers who have a greater number of common followers than expected at random. In this context, we introduce the following measures (see Materials and Methods for mathematical definitions):
(1) Community Overlap. Rather than being a unified atmosphere, the internet is split into fragmented communities where people with shared interests interact with each other. This phenomenon is defined as cyber balkanization by Van and Brynjolfsson in their work published in 2005 [7]. In this study, ordinary users can choose to join multiple communities, indicating that they are likely to have diversified interests or join just one single community, suggesting that they are more loyal to one group. We introduce the Community Overlap Index to quantify the versatility (or number) of individuals’ community membership.
(2) Identity Diversity. According to social identity theory and social categorization theory, people may act based on inter-group behaviors arising from a sense of belonging to certain social groups [35, 36]. It can be applied to the political domain, and political identities may advance social identities, such as national identities, partisan affiliations, and political ideologies. [37]. Such identity-based grouping can be extended and amplified by the digital world, as any users can reveal their identities online [38]. The diversity of influencers’ political identity in a community is measured by the Identity Diversity Index. We only consider political ideology as the main indicator of political identity (see more explanation in section 5 of the SI).
(3) Information Diversity. Instead of being exposed to certain political identity groups, ordinary users may focus on the information retrieved from specific communities. People may have a preference for certain information sources due to personal interests. Alternatively, they can be attracted by similar opinions delivered by the information. This can be explained from a psychological perspective, where people prefer to receive opinions that they agree with, a phenomenon known as confirmation bias [39, 40] or cognitive bias [41]. Here, we measure the extent to which individuals are exposed to diverse information using the website domain links posted by political influencers.
(4) Structural Isolation. Even though the online network can be segmented into numerous clusters or communities, each community has a distinctive structure. For instance, some communities are more isolated from other communities, while others are more integrated into the whole network. In the influencer network, a more isolated community indicates that it has a more targeted consumer group rather than a broader audience. We measure the extent to which individuals are exposed to isolated influencers using the Structural Isolation Index, measured by the Normalized Cut created by a community (see more details in the Methods and Materials section).
(5) Connectivity Inequality. In terms of the inner structure of influencer communities, some communities demonstrate an unequal distribution of degrees, with nodes having a larger number of connections potentially playing a more important role. We employ the Gini Index to quantify the Connectivity Inequality Index, which measures the extent to which consumers are exposed to influencer communities with more imbalanced influential power.
Having measured the patterns of consumers’ selective exposure, we continue to examine the association of ten groups of individual attributes obtained from the survey: Demographics, News Consumption, Political Communication, Political Identification, Political Engagement, Perceptions of Incivility, Perceptions of Disinformation, Authority Trust, Populism, and Attitudes toward Democracy, with the five selective exposure indices (see section 6 in SI for detailed description of the ten groups of individual attributes).
2 Results
2.1 Multidimensional Profiles of Political Influencers and Their Interconnections
We build a bipartite network between 204 political consumers and 2,307 political influencers by combining datasets of surveys and Twitter. Political consumers are sampled from respondents of a national survey conducted during the 2022 Brazilian Presidential Election. Political influencers are identified based on the Twitter accounts followed by these survey respondents and can be politicians, news media, and individual opinion leaders. Finally, 4,107 following pairs are found in the followship network. To characterize whom these consumers are following, we annotate the identities of political influencers on four dimensions: Political Ideology, Campaign Support, Social Identity, and Account Type.
Based on our annotation of political influencers on multiple dimensions, we find that influencers’ attributes across different dimensions are highly correlated (see Fig.1). A Fisher’s exact test confirms that accounts with left-wing ideologies are associated with a support of Lula - the leftist leader and current president of Brazil and advocate for women’s rights (Fisher’s Exact Test $p$-value $<0.01$). Conversely, right-wing accounts are more likely to support Bolsornaro - the rightist leader, former president of Brazil, and be associated with religious groups ($p<0.01$). Most accounts that express support for electoral candidates are from individual opinion leaders, showing only weak dependency between dimensions of Campaign Support and Account Type ($p=1.00$). However, left-wing accounts demonstrate higher proportions of politician accounts compared to others, and religious accounts show higher correlations with individual opinion leaders compared to other identities ($p<0.01$).
Figure 1: Contingency tables of multidimensional annotations of political influencer Twitter accounts. Influencer accounts are annotated based on the dimensions of Political Ideology ($N$ = 886), Campaign Support ($N$ = 356), Social Identity ($N$ = 438), and Account Type ($N$ = 2,129), with sub-categories within each dimension. Correlations among categories are checked for (a) Political Ideology vs Campaign Support, (b) Political Ideology vs Social Identity, (c) Political Ideology vs Account Type, (d) Campaign Support vs Social Identity, (e) Campaign Support vs Account Type, and (f) Account Type vs Social Identity. A multivariate Fisher’s exact test is conducted for each contingency table to examine the dependency between dimensions. All pairs of dimensions except for Campaign Support and Account Type demonstrate a strong association at 1% significance level.
2.2 Hierarchical Structure of Online Political Selective Exposure
We then project the bipartite network onto the influencer side and conduct multi-scale community detection on the influencer-projected network to measure the complexity of individuals’ selective exposure structures (see the schematic diagram in Fig.2(a) & (b) and the Material and Methods section for more details.).
In network science, community detection is used to find groups of densely connected nodes and is widely exploited in topics like selective exposure, echo chambers, and polarization. To capture the multi-scale and hierarchical organization of real-world social networks[42, 43], we conduct multi-scale scale community detection[44, 45] on the influencer network (see details in the Materials and Method and section 4 in SI). We detect five significant levels (or scales). The number of influencer communities detected at five levels is shown in Fig.2(c), ranging from 46 communities at the lowest level and 2 communities at the highest level. In this setting, a community represents a set of political influencers more likely to be followed by the same set of consumers.
Figure 2: Schematic diagram of network construction and community detection. (a) The procedure used to construct the bipartite network between political consumer nodes (orange triangles) and political influencers nodes (blue circles). We project the bipartite network onto political influencers. The directed arrows between consumers and influencers represent the following relationships on Twitter. The undirected lines between influencers indicate their shared consumers, weighted by the number of shared consumers. (b) Illustration of multi-scale community detection in the projected influencer network. At the lower level, there are more communities, and the size of each community is smaller (shown in grey). At the higher level, the smaller communities merge, resulting in fewer, larger communities (shown in orange). (c) Histograms of community sizes at the five detected levels.
Fig.3 shows the results of the multi-scale community detection of the influencer network as an alluvial diagram, revealing its hierarchical structure. As we can see, at the most granular level (Level 1), diverse communities are formed based not only on ideologies but also on political support, social group affiliation, and account types. At the second scale (Level 2), communities with mostly the same categories merge, resulting in more unified left-wing and right-wing communities, as well as distinct Lula and Bolsonaro camps. From level 3 onward, the left-wing and right-wing communities, along with Lula and Bolsonaro supporters, merge together. At the same time, many categories based on social identity and account types become less prominent. By levels 4 and 5, we observe a large mixed community and predominantly right-wing communities and Bolsonaro supporters on the side, who are more isolated and mainly come from religious groups and individual opinion leaders.
The results show that a single scale in community detection may not always detect the polarized selective exposure pattern between left-wing and right-wing groups. Lower-level scales can uncover more fragmented communities nested within ideological groups, while higher-level scales may capture more integrated ideological communities.
Figure 3: Alluvial diagram showing the hierarchical organization of political influencer communities across five levels. The different panels display the labeling of political influencer communities according to the following dimensions: Political Ideology (a), Campaign Support (b), Social Identity (c), and Account Type (d). The partition of the influencers in communities from a granular level to coarser levels is shown by the different grouping and arrangement of the bundles. We assign a color to a community if more than half of the influencer accounts within the community belong to one category. The transparency of the colors is proportional to the homogeneity of the labels; the higher the proportion of the majority category, the more salient the color. We only color communities when more than 30% of influencer accounts are annotated.
2.3 Measurements of Online Political Selective Exposure Patterns
We also project the consumer-influencer bipartite network onto the consumer side. We group consumers together if they follow political influencers in the same influencer community. The consumers can, therefore, be involved in multiple or just a single influencer community (see Fig.4(a)). We observe that as the level increases, the number of individuals involved in multiple influencer communities decreases. In contrast, the number of individuals who belong to only a single influencer community increases.
Consumers who are involved in multiple influencer communities create closer connections between these influencer communities. We visualize the influencer community network connected by shared consumers in Fig.4(b), where each pie chart represents an influencer community node and shows the proportions of different identity categories across four dimensions. The edges represent consumers who follow both communities. At Level 1 and 2, we observe that influencer communities closely connected by shared consumers exhibit high proportions of left-wings and news media categories. Consumers following these categories of influencer accounts are more versatile and fragmented than others, and this pattern can be mainly observed at the lower level.
Figure 4: Projection of the consumer-influencer bipartite network onto both the consumer and influencer sides. (a) Influencer communities projected on the consumer network. Each node represents a consumer who can follow single or multiple influencer communities. Consumers following the same communities are encircled. Consumer nodes that follow a single community are annotated in green, while those that follow more than one community are annotated in purple. All five levels are displayed. (b) Projected pie-chart graph of the influencer community network. The nodes are influencer communities, indicated by a pie chart of labels on four dimensions: Political Ideology, Campaign Support, Social Identity, and Account Type. Proportions of categories in each dimension are shown inside the pie chart. Edges between two nodes denote the consumers who follow both influencer communities.
To capture the versatility of each consumer at various levels, we develop the Community Overlap Index, which is calculated by the number of influencer communities each consumer interacts with. Indices of Identity Diversity, Information Diversity, Structural Isolation, and Connectivity Inequality are introduced to capture selective exposure patterns at the community level. Identity Diversity is measured by the diversity of political ideologies within an influencer community, and Information Diversity is measured by the diversity of website domain links shared within each community. Diversity is calculated using the Gini-Simpson Diversity Index. Additionally, we employ the Normalized Cut and Gini Index to measure Structural Isolation and Connectivity Inequality of each influencer community, respectively (see details of measurement development in the Materials and Methods section).
The values of the four community-level indices are visualized in Fig.5. The granular levels (Level 1 and Level 2) demonstrate diverse values across the four indices. As the levels increase, there is a large merged influencer community where the political ideologies are more diverse and the degree distribution is more unequal compared to other communities. Information Diversity and Structural Isolation show little distinction among communities at higher levels.
Figure 5: Visualization of measurements of four community-level indices in the influencer community network: (a) Identity Diversity, (b) Information Diversity, (c) Structural Isolation, and (d) Connectivity Inequality. The nodes in each network represent detected influencer communities, and the edges represent the shared consumers between two influencer communities. The area of each node is proportional to the number of political influencers in the community. The color spectrum ranges from black to blue, with a darker color indicating more homogeneous ideological identities and domain information, less isolation from other communities, and a more equal distribution of degree weights.
2.4 Individual Attributes Related to Multi-level Selective Exposure Patterns
We investigate how certain consumer attributes are related to selective exposure patterns at various levels with a regression analysis. Individual attributes are conceptualized into 10 groups, encompassing 189 variables, including Demographics, News Consumption, Political Communication, Political Identification, Political Engagement, Perceptions of Incivility, Perceptions of Disinformation, Authority Trust, Populism, and Attitudes toward Democracy. We conduct Principal Component Analysis (PCA) within each variable group to address the issue of multicollinearity. This results in 25 variables after PCA dimension reduction, which are then regressed against the five indices. It is worth noting that we convert the four community-level indices — Identity Diversity, Information Diversity, Structural Isolation, and Connectivity Inequality — into individual-level indices by calculating the average values of the influencer communities each consumer interacts with, weighted by the number of influencers each consumer follows.
We identify significant explanatory variables among the 25 initial variables for each index at each level using a forward selection approach based on the Bayesian Information Criterion (BIC). This model selection procedure aims to find the statistical model with the fewest explanatory variables by iteratively adding them until the statistical significance of the model, as measured by the BIC, no longer improves (see Material and Methods for details).
We fit the minimal explanatory variables into the regression model for each index and each level. After conducting distribution checks for the values of dependent variables, the zero-truncated Negative Binomial Regression model is chosen as the basic model for the Community Overlap Index. For the indices of Identity Diversity, Information Diversity, Structural Isolation, and Connectivity Inequality, the Beta Regression model is employed. For more details, refer to the Materials and Methods and to section 6 in SI. The regression results for the first four levels are shown in Fig.6.
For the Community Overlap Index, we can observe that the ideological position of individuals plays a significant role in explaining their involvement in multiple communities at all levels. However, it is not the only factor; at lower levels of communities, more factors contribute to the versatile engagement of consumers in influencer communities (see Fig.6(a)). The impact of ideological position corresponds to the fragmentation pattern displayed in Fig.4(b), suggesting that left-wing influencer communities are more likely to be shared by versatile consumers. Another two significant factors that remain at the first two levels are the frequency of news consumption on Twitter and age. At Level 3, the sign of the ideological position coefficient changes, demonstrating a turning point from which right-wing consumers are more inclined to be involved in multiple influencer communities. This is because as the left-wing influencers merge, most of the remaining isolated communities are right-wings. The ideological position plays an increasingly important role at higher levels, especially at Level 4, where only the ideological position remains significant.
The other community-level indices show correlations with different attributes, but similarly to the Community Overlap, they demonstrate various associations across levels. For instance, Military Trust primarily explains Identity Diversity at Level 1. At Level 2, the frequency of individuals participating in political activities is a more prominent factor (see Fig.6(b)). Demographic variables play a more important role at Levels 3 and 4, where the religious level is consistent at both levels.
The ideological position of individuals plays a more significant role in Information Diversity compared to other indices. As shown in Fig.6(c), the ideological position is the only significant factor at Levels 1, 3, and 4, and a sub-dimension of populist ideology — people centralism, which is the demand for prioritizing the rights of the majority, is a significant factor at Level 2.
For Structural Isolation, there are many explanatory factors at the finer levels, while the degree of isolation of communities is mainly related to consumers’ support for democracy at Levels 3 and 4 (see Fig.6(d)). Regarding Connectivity Inequality, similar to Community Overlap, at lower scales, ideological position, as well as a variety of other factors such as age, gender, religious level, and the frequency of news consumption on Twitter, have explanatory power. Only the ideological position remains significant at Level 4 (see Fig.6(e)).
Figure 6: Regression plot of significant individual attributes related to selective exposure patterns. Regression analyses of consumers’ attributes on Demographics, News Consumption, Political Communication, Political Identification, Political Engagement, Perceptions of Incivility, Perceptions of Disinformation, Authority Trust, Populism, and Attitudes toward Democracy are conducted across five indices of selective exposure patterns: (a) Community Overlap, (b) Identity Diversity, (c) Information Diversity, (d) Structural Isolation, and (e) Connectivity Inequality at four network community levels. The variables displayed are significant factors resulting from a forward selection based on BIC. The coefficient values are displayed in blue, and the confidence intervals at the 95% confidence level are shown with error bars. The red lines indicate the 00 value.
3 Discussion and Conclusion
The findings of this study reveal the intricate and hierarchical nature of individuals’ political selective exposure in the Brazilian Twitter space. By employing a multidimensional annotation design, we discover that attributes of political influencers are highly correlated across different dimensions. For instance, left-wing accounts are correlated with attributes of supporting Lula and advocating for women’s rights. In contrast, right-wing accounts are more likely to be associated with Bolsonaro supporters and religious believers. This highlights political influencers involved in homogeneous left-wing and right-wing communities or Lula-camp and Bolsonaro-camp also reassemble each other on demographic grounds. Such granularity is important because it suggests politicians and political parties may employ more targeted mobilization strategies to gain support from societal subgroups beyond the traditional ideological camps.
In the long history of political studies, demographics such as gender, race, ethnicity, and religion have been related to ideological positions and voting behaviors [46]. However, in recent years, demographic cleavages have become more significant due to the emergence of far-right and populist political leaders in political campaigns [47]. Layton et al. (2021) demonstrated a pattern of new alignments in Brazil’s electorate since the 2018 presidential election, showing that political groups are now split more based on demographics and issues rather than classic partisanship [34]. In particular, the far-right candidate Bolsonaro has attracted new supporters by creating demographic polarization beyond the traditional left-versus-right division. For instance, his misogynistic views on gender, racially offensive remarks to Black activists, and evangelical propositions have played a role in this polarization [34]. The strong correlations between political ideology and other dimensions suggest that a new trend of political division might be ongoing.
The hierarchical patterns observed through multi-scale community detection demonstrate the importance of levels in understanding political selective exposure. At the most granular level, diverse communities form based on various factors, including political ideology, candidate support, social group affiliation, and account type. As the scale increases, these communities gradually merge, forming more unified ideological groups. This process continues until higher scales predominantly reveal right-wing communities and Bolsonaro supporters as homogeneous and isolated communities, which are mainly associated with religious groups and individual opinion leaders. These results show that single-scale community detection methods may miss the nuanced, nested structures within political communities, underscoring the value of multi-scale approaches in capturing both fragmentation and unification patterns.
Our investigation into the related factors of selective exposure patterns reveals that different factors play a role in multi-level selective exposure. We find that Community Overlap is a useful measure of a consumer’s multi-level versatility. Many variables, such as demographics, news consumption frequency, and ideological position, are found to be related to individuals’ versatility in connecting to influencer communities. However, only political ideology is significant at the most coarse level. These findings indicate that Brazilian online users’ selective exposure is indeed influenced by non-ideological factors such as demographics, even though ideological division simultaneously exists at multiple levels. We should note that in the long run, such divisions could be enlarged and threaten the stability of society. Scholarly, it also suggests that online communities detected at a single level might miss important information; therefore, various levels should be examined in relevant studies.
This study builds both theoretical and methodological foundations for exploring hierarchical selective exposure and related concepts such as political fragmentation, polarization, and echo chambers. Future studies could explore more countries and platforms and also examine the dynamics over time. Additionally, other identities of political influencers could be labeled, and alternative indices of selective exposure patterns could be tested. In summary, this study demonstrates the necessity of employing multi-level approaches to effectively capture the complexity of political selective exposure on social media platforms like Twitter. Our results highlight that political communities are not monolithic; rather, they are characterized by intricate interconnections across various dimensions and scales. Single-scale analyses may overlook critical aspects of community formation and evolution, potentially leading to incomplete or biased understandings of online political patterns. Future research should continue to refine such multi-scale methodology, potentially combining with content and issues, to interpret the multi-level behaviors of minority groups and radical ideological groups.
4 Materials and Methods
4.1 Data Collection and Preprocessing
@@
During the 2022 Brazilian presidential election (from October to November 2022), a national survey was conducted by NetQuest, an international survey company. The survey involved 1,018 respondents who were representatively sampled in terms of gender and region. Additionally, the respondents met a targeted income distribution across the nation. Approximately 40 questions and 189 variables regarding socio-demographics, news consumption behaviors, and political predispositions were asked in the survey (see the questionnaire in section 1 in the SI). We also requested respondents’ consent to provide their Twitter handles. Of the 1,018 respondents, 403 consented, and 271 were verified as existing accounts on Twitter, accounting for 26.62% of the whole panel. A Kolmogorov-Smirnov (K-S) test of the demographic variables between the 271 samples and the 1,018 respondents does not reject the null hypothesis that both samples are drawn from the same distribution at a significance level of 5% (see Fig. S1 in SI). We therefore consider this 271-subsample statistically representative of the entire group of respondents.
@@
Based on the 271 survey respondents, we collected the Twitter accounts followed by them (57,645 accounts and 73,755 following pairs were collected) using the Twitter API. We use “following” relations to represent individuals’ preference to be exposed to information posted by the accounts they follow. We then identify political influencers from the 57,645 accounts. We define political influencers as a composition of both ordinary citizens and celebrities (e.g., politicians, parties, media outlets, journalists, and individual opinion leaders) who satisfy two conditions: 1) with at least more 1,000 followers and 2) are likely to produce political content [49, 50] (see section 2 in SI for more details about the identification of political influencers). Finally, we obtained 204 survey respondents (referred to as political consumers in the text), 2,307 political influencers, and 4,107 following pairs between the two groups.
4.2 Multi-scale Community Detection

### v2

# A multilevel network approach to revealing patterns of online political selective exposure

@@ Abstract
Selective exposure, individuals’ inclination to seek out information that supports their beliefs while avoiding information that contradicts them, plays an important role in the emergence of polarization and echo chambers. In the political domain, selective exposure is usually measured on a left-right ideology scale, ignoring finer details. To bridge the gap, this work introduces a multilevel analysis framework based on a multi-scale community detection approach. To test this approach, we combine survey and Twitter/X data collected during the 2022 Brazilian Presidential Election and investigate selective exposure patterns among survey respondents in their choices of whom to follow. We construct a bipartite network connecting survey respondents with political influencers and project it onto the influencer nodes. Applying multi-scale community detection to this projection uncovers a hierarchical clustering of political influencers, where each cluster is more frequently co-followed by a specific subgroup of survey respondents compared to others. Different indices of selective exposure, such as Community Overlap, Identity Diversity, Information Diversity, Structural Integration, and Connectivity Inequality, suggest that the characteristics of the influencer communities engaged by survey respondents vary with the level of community resolution. This finding indicates that online political selective exposure exhibits a more complex structure than a mere left-right dichotomy. Moreover, depending on the resolution level we consider, we find different associations between network indices of exposure patterns and 189 individual attributes of the survey respondents. For example, at finer levels, survey respondents’ Community Overlap is associated with several factors, such as ideological position, demographics, news consumption frequency, and incivility perception. In comparison, only their ideological position is a significant factor at coarser levels. Our work demonstrates that measuring selective exposure at a single level, such as left and right, misses important information necessary to capture this phenomenon correctly.
Keywords: Political Selective Exposure $|$ Political Polarization $|$ Echo-Chambers $|$ Social Media $|$ Social Network $|$ Multi-scale Community Detection
Introduction
@@
The internet, especially social media, has facilitated the online consumption of political information in recent decades [1]. According to a national survey we conducted during the 2022 Brazilian Presidential Election, 82 percent of the Brazilian population acquired political information via social media, with approximately 51 percent of them consuming or engaging with political information on Twitter (now known as X) at least once a week. On Twitter/X, ordinary users mainly receive political information by following certain political influencers, such as politicians, media outlets, journalists, and individual opinion leaders. Despite social networks like Twitter/X providing a high-choice environment for information consumption and the possibility of deliberative communication, it is a concern that users on social networks predominantly follow information aligned with their own opinions, leading to political fragmentation, polarization, and the creation of echo chambers [2, 3, 4, 5, 6]. This phenomenon, known as selective exposure, is attributed to bounded rationality—humans have limited capacity for information processing and prefer to be exposed to information they find agreeable [7] and to the influence of personalized recommendation algorithms on social media platforms [8, 9, 10, 11, 12, 13]. Selective exposure online can, therefore, limit people’s access to diverse opinions and even lead to the spread of extreme views within certain groups [14].
Many studies have explored selective exposure and related topics such as political polarization and echo chambers, e.g., [15, 16, 17, 18, 19, 20, 5, 21, 22]. Most of them focus on the U.S. case, where the political system has traditionally been divided between two major parties - the Democrats and the Republicans [23]. A similar left–right division is evident in users’ behaviorally selective exposure, where individuals are more likely to interact with members of their own party [24]. This phenomenon deepens polarization both ideologically and behaviorally through the reinforcement of echo chambers [25]. However, the dichotomous model of selective exposure, polarization, or echo chambers might be not fully applicable to countries like Brazil, which has a multi-party system and weak party affiliation. Scholars often observe non-ideological cleavage fragmentation in multi-party systems, although the left–right ideological divide persists [26]. Therefore, this study focuses on Brazil as a case study, where both non-ideological fragmentation and ideological division may contribute to the complex patterns of selective exposure among Brazilian online users.
@@
Furthermore, user-generated content on social media has reshaped the dichotomous nature of online political dynamics globally[27, 5, 29]. It is not only elite accounts such as media outlets, political parties, and politicians that promote political propaganda, but also individual opinion leaders who create more targeted political information to represent specific social groups or achieve business attention [30, 31]. In light of this background, political division is more complex in the contemporary era than ever. Evidence from the 2016 U.S. presidential election presented by Bartels et al. [33] highlights that within the Democratic group, various cultural subgroups exist where social identities and other demographic characteristics play a role in forming closely connected communities. Similarly, Brazil’s party systems are also intertwined with complex socio-cultural groups that could even transcend party lines [32].
That said, traditional network community detection algorithms that perform well in identifying left–right selective exposure, polarization, and echo chambers in the U.S. are often inadequate for uncovering fragmentation patterns in countries like Brazil or in the current social media age. Modularity-based quality functions suffer from a resolution limit, making them ineffective at detecting smaller communities within large networks and at multiple resolution levels [34]. Hence, in this study, we apply a multi-scale community detection framework based on Markov stability to uncover patterns of political selective exposure among Brazilian online users sampled from a representative national survey across varying levels of resolution [35]. While other network-based approaches can also measure polarization in multipolar online systems, our approach offers a distinct advantage in uncovering the hierarchical structure of political systems [61, 62]. Additionally, we propose five novel indices of selective exposure and examine their associated factors at various resolution levels. For each surveyed individual, they are computed based on the communities that their followed influencers form. They are: (i) Community Overlap, capturing the bridging capacity of the surveyed individuals; (ii) Identity Diversity, measuring the diversity of the identities of influencers that surveyed individuals follow; (iii) Information Diversity, capturing the information provision diversity of their influencers; (iv) Structural Integration, measuring the network integration of their influencers; and (v) Connectivity Inequality, capturing the influence imbalance of their influencers. Overall, our research is guided by two main questions:
  • RQ1:
    How can we capture the complexity of individuals’ selective exposure, reflected in Community Overlap, Identity Diversity, Information Diversity, Structural Integration, and Connectivity Inequality, within a fragmented political context?
  • RQ2:
    Which individual attributes are associated with selective exposure patterns across different resolution levels?
@@
To answer the two research questions, we combine a representative national survey and Twitter/X political influencers the survey participants followed, and construct consumer-supplier relations. Consumers are represented by 204 respondents from a national survey who provided their Twitter/X handles, while the suppliers are 2,307 political influencers (including politicians, media, and individual influencers) followed by consumers. The distribution of demographic attributes of the 204 survey participants does not show significant statistical differences compared to the overall survey sample, which is representatively sampled from Brazil’s population. Our analysis, therefore, reports on selective exposure from the point of view of a representative sample of users. Moreover, political influencers are annotated based on dimensions of political ideology (Left/Right/Center), campaign support (Lula camp/Bolsonaro camp), social identity (Women/Religious/LGBTQ/Black), and account type (Politician/Media/Individual). Then, we uncover the complex structures of Brazilian users’ selective exposure on Twitter/X using a multi-scale community detection approach. The results suggest that ordinary users tend to expose themselves to political influencers based on ideological positions, however, selective exposure patterns are also evident along other dimensions, such as electoral supports, social identities, and categories of accounts. Therefore, Brazilian users’ selective exposure on Twitter/X forms a multi-level and hierarchical structure.
@@
By examining a comprehensive set of individuals’ survey attributes with the five selective exposure indices at various levels, we find that, at the lowest level, a wide range of variables, including ideological position, demographics, news consumption frequency, and incivility perception, are significantly associated with the survey individuals’ Community Overlap. As the level increases, the ideological position becomes more prominent. The other indices are also associated with different individual attributes at various levels. These findings suggest that the multilevel framework based on multi-scale community detection is an effective approach for revealing individuals’ selective exposure across different levels of resolution, and it can be also applied to other contexts and larger datasets.
@@
This study significantly contributes to the application of multi-scale community detection in social network research, demonstrating that different resolution levels can reveal distinct patterns of individuals’ selective exposure. This approach is particularly valuable for studying selective exposure, polarization, and echo chambers in multi-party political systems or societies with complex social structures.
Materials and methods
Data collection and preprocessing
@@
During the 2022 Brazilian presidential election (from October to November 2022), a national survey was conducted by NetQuest, an international survey company. The data were gathered using NetQuest’s exclusive online panel, using stratified quota sampling aligned with national distributions of age, gender, and geographic area to ensure national demographic representativeness. Statistical tests indicate that the survey participants constitute a representative sample of the national electorate (see Fig B in Supplementary Information). Participants were enrolled through a double opt-in process and completed the survey online in return for incentives. Approximately 40 questions and 189 variables regarding socio-demographics, news consumption behaviors, and political predispositions were asked in the survey (see the questionnaire in section 2 in Supplementary Information). We also requested respondents’ consent to provide their Twitter/X handles. Of the 1,018 respondents, 403 consented, and 271 were verified as existing accounts on Twitter/X, accounting for 26.62% of the whole panel.
@@
Based on the 271 survey respondents, we collected the Twitter/X accounts followed by them (57,645 accounts and 73,755 following pairs were collected) using the Twitter/X API. We use “following” relations to represent individuals’ preference to be exposed to information posted by the accounts they follow. We then identify political influencers from the 57,645 accounts. We define political influencers as a composition of both ordinary citizens and celebrities (e.g., politicians, parties, media outlets, journalists, and individual opinion leaders) who satisfy two conditions: 1) with at least more 1,000 followers and 2) indicate political related information in their profiles or are known political figures [1, 2] (see section 3 in Supplementary Information for more details about the identification of political influencers). After filtering out the survey respondents who did not follow any political influencers, we obtained 204 survey respondents (also referred to as political consumers in the text), 2,307 political influencers (also referred to as political suppliers in the text), and 4,107 following pairs between the two groups. Pearson’s chi-squared tests (for categorical variables) and Mann–Whitney U tests (for discrete variables) comparing the demographic distributions of the 204 subsample with the 1,018 survey respondents do not reject the null hypothesis that the 204 samples are drawn from the same distribution as the survey respondents at the 5% significance level (see Fig C in Supplementary Information). Therefore, the survey respondents who provided their handles are very likely to form a representative sample of the Brazilian electorate.
This study was reviewed and approved by the Ethics Committee of the Faculty of Arts and Social Sciences, University of Zurich. Informed electronic consent was obtained prior to participation via the NetQuest survey platform. Participants were informed of the study purpose, data processing, voluntary participation, and their right to withdraw at any time. Twitter/X data were collected in accordance with the platform’s Terms of Service and API policies and were limited to publicly available content.
Multi-scale community detection